OpenStudy (anonymous):

On a particular day, the wind added 3 km per hour to Alfonso’s rate when he was cycling with the wind and subtracted 3 km per hour from his rate on his return trip. Alfonso found that in the same amount of time he could cycle 36.5 km with the wind, he could go only 24 km against the wind. What is his normal bicycling speed with no wind? a. 15 km/h c. 35 km/h b. 40 km/h d. 45 km/h

OpenStudy (anonymous):

the general equation for this is Distance = Rate x Time Let Alfonso's normal biking speed = R and let T be the length of time in the problem 36.5=(R+3) x T 24=(R-3) x T solve for R

OpenStudy (aravindg):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school, I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. Welcome to the OpenStudy Community! Ambassador

OpenStudy (anonymous):

As the question stands, none of the multiple choice answers are correct. Answer A is close, but the answer is not exactly a nice number. For answer A to be correct, instead of being able to cycle 36.5 km in the length of time, Alfonso should only be able to cycle 36 km.

OpenStudy (anonymous):

@Peter14 I dont know how to solve further for R in the equations mentioned at the top

OpenStudy (anonymous):

@ganeshie8 @mathmate

OpenStudy (mathmate):

@peter14 has already set up the equations: 36.5=(R+3) x T 24=(R-3) x T You will need to solve for R. To do this, you would divide one equation by the other to eliminate T. 36.5/24 = (R+3)/(R-3) Cross multiply to get 36.5(R-3)=24(R+3) Expand and solve for R. As Peter said, it will be a not nice number, but close to 15 km/h. There is probably a typo in either the question or the answers.

OpenStudy (anonymous):

okay thanks

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

hahaaa maimi girl..... coincidently @jaikid_077 and i had the same question

OpenStudy (mathmate):

ok, that explains it! :)

OpenStudy (anonymous):

good question though...lol

OpenStudy (mathmate):

:) talk to you some other time!